OpenStudy (anonymous):

Can someone help me understand this please? I'm not asking for you to do it for me just show me how. A = batting average B = times at bat multiplied by player’s seasonal average K = hits in the new game G = times at bat plus times at bat in a new game Solve A = (b+k) over g for K. Please show your work. Use your new formula and the values given in the table below to find the player’s hits in a new game (K). Round your answer to the nearest whole number

OpenStudy (anonymous):

Batter A B G K George 0.319 10.7 46 Jose 0.320 11.4 45 Dale 0.298 12.3 52 Michael 0.337 15.2 48

OpenStudy (anonymous):

Average A = ( B - K ) / G -> AG = B - K Thus the player hit in the new game K = B - AG Now plug the date from table: George: = 10.7 - ( .319 * 46 ) = ....